psirockin2:

Write an equation of a circle with the given center and radius center (3,4) and radius 6

pandasurvive:

1 attachment
Nnesha:

\[\huge\rm (x-h)^2+(y-k)^2=r^2\] where (h,k) is the center point and r is the radius

psirockin2:

um Nnesha, is that LaTex you're using? I can't see it (K9 started blocking Latex too)

Nnesha:

(x-h)^2+(y-k)^2=r^2

Elsa213:

`[Math Processing Error]`

psirockin2:

Okay so I stick the numbers in that formula Nnesha?

Nnesha:

7b5b4a7630a14056a19340a3ad981fb4.png

Nnesha:

yes

psirockin2:

Okay I think I can do it. I always forget the formulas for each thing since there's like 6 they want us to memorize

Nnesha:

Practice also helpz. If you solve at least 10 same type of questions you will be fine!

psirockin2:

okay thanks!

Nnesha:

yw :=))